OpenStudy (anonymous):

Integrate 5dx / Sqrt (16-25x^2)?

OpenStudy (anonymous):

\[\sin^{-1} (\frac{ 5 }{ 4 }x)\]

OpenStudy (anonymous):

I see.. can you show me steps on how to do it?

OpenStudy (anonymous):

\[\frac{ 5 }{ \sqrt{16-25^{2}} } \] \[\frac{ 5 }{ 16(1-\frac{ 25^2 }{ 16 } }\] \[\frac{ 5 }{4(\sqrt( 1-(\frac{ 5x }{ 4 })^2 }\]

OpenStudy (anonymous):

are you able to see it?

OpenStudy (anonymous):

Yes...Thanks so much

OpenStudy (anonymous):

okay.. your welcome